OpenStudy (jakob):

If the range of the function f(x)= 5x − 3 is {-13, -8, -3, 12, 17}, what is its domain? A{0, 1, 2, 3, 4} B{-2, -1, 0, 1, 2} C{-2, -1, 0, 3, 4} D{-4, -3, 0, 3, 4}

OpenStudy (welshfella):

the range of 5x - 3 is the possible values of 5x - 3 the domain is the corresponding values of x For example if 5x - 3 = -3 what is the value of x?

OpenStudy (welshfella):

solve this equation for x and you have one element of the domain.

OpenStudy (jakob):

I am lost

OpenStudy (welshfella):

5x - 3 = -3 5x - 3 + 3 = -3 + 3 5x = 0 so x = what?

OpenStudy (welshfella):

the possible values in the range are given to you in the set range = {-13 , -8, -3 , 12, 17}

OpenStudy (jakob):

Idk what x is

OpenStudy (welshfella):

I just took the member -3 and equated it to 5x - 3 in order to find a member of the domain (x)

OpenStudy (maddielb):

Where did you get the second -3

OpenStudy (welshfella):

5x = 0 5 * 0 = 0 so x = 0

OpenStudy (welshfella):

@maddielb : - 3 is one element on the range

OpenStudy (jakob):

Please hurry need to get this done

OpenStudy (welshfella):

so 0 is one member of the domain to find the other members you need to solve 5x - 3 = -13 5x - 3 = -8 5x - 3 = 12 and 5x - 3 = 17